# Combinatorics and Hodge theory of degenerations of abelian varieties: A survey of the Mumford construction

### v2

@@ 6. The second Voronoi fan and Alexeev’s theorem
Let the bilinear forms $B_{1},\dots,B_{k}$ be extremal rays of a polyhedral monodromy cone ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$ and let $X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ be a universal polytopal Mumford Construction 3.25, resp. let $X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ be a fan Mumford Construction 3.10. Then, $X^{\rm univ}(\mathbbm{b})$, resp. $X^{\rm univ}({\mathcal{S}})$, arise by restricting to an analytic open subset of the base, a projective morphism of quasiprojective varieties, resp. algebraic morphism of algebraic spaces.
@@
By replacing $\widetilde{{\mathcal{A}}}_{g}$ with a suitable further cover, we may assume that the disinction between ${\mathcal{X}}_{g}$ and ${\mathcal{X}}_{g}^{\star}$ is erased. Let $\mathfrak{F}$} be a common refinement of the fans $\Gamma\cdot{\mathbb{B}}$ and ${\mathfrak{F}}_{\rm vor}$ whose support is $\Gamma\cdot{\mathbb{B}}$, for $\Gamma\subset\operatorname{GL}_{g}({\mathbb{Z}})$ the Levi quotient. Then we have morphisms
@@
and we may pullback the (a priori) analytic, relatively projective family $X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ and the algebraic universal family $\widetilde{{\mathcal{X}}}_{g}^{\rm vor}\to\widetilde{{\mathcal{A}}}_{g}^{\rm vor}$ to produce two families $X^{\rm univ}\,\!{}^{\prime}(\mathbbm{b})$, $X^{\rm vor}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathfrak{F}}}$ in the analytic and algebraic categories, respectively.
@@
We now observe that by taking a common refinement of the polyhedral decompositions ${\rm Del}_{B}$ and $\bigcup_{i=1}^{k}{\rm Bend}(b_{i})$ of ${\mathbb{T}}^{g}$ (or in the case of the fan construction, a common refinement of the relevant fans), we may dominate $X^{\rm univ}\,\!{}^{\prime}(\mathbbm{b})$ and $X^{\rm vor}$ by a common (universal) Mumford construction ${\widetilde{X}}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathfrak{F}}}$. Finally, ${\widetilde{X}}\to X^{\rm univ}\,\!{}^{\prime}(\mathbbm{b})$ and ${\widetilde{X}}\to X^{\rm vor}$ are both toroidal morphisms, and hence it is possible to connect $X^{\rm vor}\dashrightarrow X^{\rm univ}\,\!{}^{\prime}(\mathbbm{b})$ by a sequence of blow-ups and blow-downs of toric strata, all of which are algebraic. The proposition follows—the relative projectivity/algebraicity of $X^{\rm univ}\,\!{}^{\prime}(\mathbbm{b})$ and $X^{\rm univ}(\mathbbm{b})$ are equivalent, since $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathfrak{F}}}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ is also a proper algebraic modification. ∎
@@
The goal of this section is to prove Theorem 1.1, and leverage Proposition 6.12 to prove more algebraic formulations of the results therein, see Theorem 7.1 and Corollary 7.2 below.
@@
Let $f^{\ast}\colon X^{\ast}\to(\Delta^{*})^{k}$ be a family of PPAVs which is matroidal with respect to the snc extension $(\Delta^{*})^{k}\hookrightarrow\Delta^{k}$ (Def. 4.9). Then there are integers $r_{i}>0$ for which the monodromy bilinear forms about $\{u_{i}=0\}$ are $r_{i}B_{i}$ where $B_{i}={\bf x}_{i}^{2}$ is an integral generator of the matroidal cone ${\mathbb{B}}_{\underline{R}}$ of the corresponding regular matroid $\underline{R}$}. Constructions 4.16 and 3.25 give a universal Mumford degeneration
@@
The monodromies about the boundary divisors of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$ are exactly $r_{i}B_{i}$ and by Proposition 2.27, the classifying morphism $(\Delta^{*})^{k}\to{\mathcal{A}}_{g}$ will (lift and) extend to $\Delta^{k}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$. Pulling back $X^{\rm univ}({\underline{R}},\mathscr{H})$ along the extension of the classifying morphism to $\Delta^{k}$ produces the desired extension $f\colon X\to\Delta^{k}$. It has smooth total space and nodal singularities, because $f^{\rm univ}$ is locally trivial along the deepest toroidal stratum of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$ and so smoothness, resp. nodality, of $X^{\rm univ}$, resp. $f^{\rm univ}$, implies smoothness, resp. nodality, of the restrictions $X$, resp. $f$, to the transversal slice $\Delta^{k}$ to this deepest toroidal stratum.
@@
The second part of Theorem 1.1 follows from the existence of algebraic, transversely shifted matroidal degenerations with specified monodromies, see Theorem 7.2 below. ∎
@@
Let $f^{\ast}\colon X^{\ast}\to Y^{\ast}$ be a projective family of PPAVs over a base $Y^{*}=Y\setminus D$ for $D\subset Y$ an snc divisor in a smooth quasiprojective variety $Y$. Let $0\in D$ and assume the local monodromy bilinear forms $B_{i}$ about the components $D_{i}\ni 0$ are $r_{i}{\bf x}_{i}^{2}$ for an integral realization $i\mapsto{\bf x}_{i}\in{\bf M}^{\vee}$ of a regular matroid $\underline{R}$}, where ${\bf M}\simeq{\rm gr}^{W}_{0}H_{1}(X_{t},{\mathbb{Z}})$ for $t$ near 0. Up to passing to an étale neighborhood of $0\in Y$, there is a flat, projective, relatively $K$-trivial extension $f\colon X\to Y$ of quasiprojective varieties, which is $D$-nodal, and strictly $D$-nodal if all $r_{i}\geq 2$.

### v3

@@ 6.2. Algebraicity and projectivity
Let ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$ be a rational polyhedral cone and $\mathcal{S}$} be a fan satisfying the hypotheses of Construction 3.11. The corresponding Mumford construction $f\colon X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ is a proper, equidimensional morphism of algebraic spaces.
Suppose, furthermore, that $f\colon X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ is a polytopal Mumford construction, as in Construction 3.26. Then $f$ is étale-locally projective.
@@
By replacing $\widetilde{{\mathcal{A}}}_{g}$ with a suitable further cover, we may assume that the distinction between ${\mathcal{X}}_{g}$ and ${\mathcal{X}}_{g}^{\star}$ is erased. Let $\mathfrak{F}$} be a common refinement of the fans $\Gamma\cdot{\mathbb{B}}$ and ${\mathfrak{F}}_{\rm vor}$ whose support is $\Gamma\cdot{\mathbb{B}}$, for $\Gamma\subset\operatorname{GL}_{g}({\mathbb{Z}})$ the Levi quotient. Then we have morphisms
@@
and we may pullback the (a priori) analytic family $X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ and the algebraic universal family $\widetilde{{\mathcal{X}}}_{g}^{\rm vor}\to\widetilde{{\mathcal{A}}}_{g}^{\rm vor}$ (see Construction 6.6) to produce two families $X^{\rm univ}\,\!{}^{\prime}({\mathcal{S}})$ and $X^{\rm vor}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathfrak{F}}}$ in the analytic and algebraic categories, respectively.
@@
Taking a common refinement of the fans defining $X^{\rm univ}\,\!{}^{\prime}({\mathcal{S}})$ and $X^{\rm vor}$ we may dominate $X^{\rm univ}\,\!{}^{\prime}({\mathcal{S}})$ and $X^{\rm vor}$ by a common (universal) Mumford construction ${\widetilde{X}}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathfrak{F}}}$. Since ${\widetilde{X}}\to X^{\rm univ}\,\!{}^{\prime}({\mathcal{S}})$ and ${\widetilde{X}}\to X^{\rm vor}$ are both toroidal morphisms, we can connect $X^{\rm vor}\dashrightarrow X^{\rm univ}\,\!{}^{\prime}({\mathcal{S}})$ by a sequence of toric modifications, all of which are algebraic. We deduce that $X^{\rm univ}\,\!{}^{\prime}({\mathcal{S}})$ is an algebraic space. In turn, its contraction $X^{\rm univ}({\mathcal{S}})$ is a proper algebraic space over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$.
Finally, we address the case of a polytopal Mumford construction $f\colon X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. Then $f$ is a morphism of algebraic spaces by what we have just proved, and by (a generalization of) Theorem 3.29 that compares the polytope and fan Mumford constructions over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. Moreover, $f$ is analytically-locally projective over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$, see Construction 3.26. The second statement of the proposition is therefore a consequence of the following general fact: If a separated and finitely presented morphism of algebraic spaces $X\to Y$ is analytically-locally projective, then it is étale-locally projective.
The proof follows from Artin approximation. Indeed, the Hilbert scheme $\operatorname{Hilb}_{X/Y}$ is an algebraic space locally of finite presentation over $Y$ by [63, Tag 0D01]. Take a point $p\in Y$. An analytic family of ample divisors $D_{U}\subset X_{U}\to U$ over an analytic neighborhood $U\ni p$, may be approximated by an algebraic family of ample divisors, $D^{\prime}_{U^{\prime}}\subset X_{U^{\prime}}\to U^{\prime}$ over an étale neighborhood $U^{\prime}\ni p$, which coincides with the restriction of $D_{U}$ to $p$. Possibly replacing $U^{\prime}$ with a smaller, Zariski open neighborhood of $p\in U^{\prime}$, the divisor $D_{U^{\prime}}^{\prime}$ is relatively ample. ∎
We now consider the much subtler question of when $X^{\rm univ}(\mathbbm{b})$ is projective over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$, as opposed to étale-locally projective.
Definition 6.13.
Let $\overline{b}$, $\overline{b}\,\!^{\prime}\in H^{0}({\mathbb{T}}^{g},\tfrac{1}{d}{\mathbb{Z}}{\rm PL}/\tfrac{1}{d}{\mathbb{Z}}{\rm L})$. We say that $\overline{b}\sim\overline{b}\,\!^{\prime}$ lie in the same shift class if $\overline{b}-\overline{b}\,\!^{\prime}$ lifts to an $\bf M$}-periodic section $b-b^{\prime}\colon{\bf M}_{\mathbb{R}}\to{\mathbb{R}}$ of $\tfrac{1}{d}{\mathbb{Z}}{\rm PL}$.
Recall that ${\bf M}\simeq{\mathbb{Z}}^{g}=H_{1}({\mathbb{T}}^{g},{\mathbb{Z}})$ in Construction 3.38. A necessary, but insufficient, condition for $\overline{b}\sim\overline{b}\,\!^{\prime}$ is that they define the same monodromy bilinear form $B$ via formula (25).
Example 6.14.
Let $g=1$ and consider $\overline{b}$, $\overline{b}\,\!^{\prime}$, $\overline{b}\,\!^{\prime\prime}$ for which ${\rm Bend}(\overline{b})=2[\tfrac{0}{1}]$, ${\rm Bend}(\overline{b}\,\!^{\prime})=[\tfrac{0}{2}]+[\tfrac{1}{2}]$, and ${\rm Bend}(\overline{b}\,\!^{\prime\prime})=[\tfrac{1}{3}]+[\tfrac{2}{3}]$ as $\mathbb{Z}$}-weighted linear combinations of $\tfrac{1}{d}$-integral codimension $1$ polytopes, see Definition 3.17. All three define the same monodromy bilinear form $B=2x^{2}$. But we have $\overline{b}\sim\overline{b}\,\!^{\prime\prime}$ and $\overline{b}\not\sim\overline{b}^{\prime}$. See Figure 23. The fundamental issue is that, while we could subtract from $b-b^{\prime}$ a linear function of slope $\tfrac{1}{2}$ to make it periodic, such a function is not a section of $\tfrac{1}{2}{\mathbb{Z}}{\rm L}$ on $\bf M$$_{\mathbb{R}}\simeq{\mathbb{R}}$, see Definition 3.25, because it does not have integral slope.
Figure 23. Left: lift of section $\overline{b}-\overline{b}\,\!^{\prime}$ with bending $\tfrac{0}{2}]-[\tfrac{1}{2}]$. Right: lift of section $\overline{b}-\overline{b}\,\!^{\prime\prime}$ with bending $2[\tfrac{0}{3}]-[\tfrac{1}{3}]-[\tfrac{2}{3}]$.
Proposition 6.15.
Let $f\colon X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ and $f^{\prime}\colon X^{\rm univ}(\mathbbm{b}^{\prime})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ be two universal polytopal Mumford Constructions 3.26, for lifts of two cones $\mathbbm{b},\mathbbm{b}^{\prime}\subset H^{0}({\mathbb{T}}^{g},\tfrac{1}{d}{\mathbb{Z}}{\rm PL}/\tfrac{1}{d}{\mathbb{Z}}{\rm L})$ into $H^{0}({\bf M}_{\mathbb{R}},\tfrac{1}{d}{\mathbb{Z}}{\rm PL})$, mapping to the same monodromy cone ${\mathbb{B}}\subset{\rm Sym}^{2}{\bf M}^{\vee}$. Suppose, that for any $B\in{\mathbb{B}}$, the PL lifts $b,b^{\prime}\colon{\bf M}_{\mathbb{R}}\to{\mathbb{R}}$ satisfy the property that $b^{\prime}-b$ is $\bf M$}-periodic. Then, the canonical analytic polarization on the Mumford construction extends as a relatively ample global section of the relative Picard ${\rm Pic}_{X^{\rm univ}(\mathbbm{b})/\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}}$ if and only if the same holds for $X^{\rm univ}(\mathbbm{b}^{\prime})/\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$.
We note that the data of the gluing of the Mumford construction onto the universal family $\widetilde{{\mathcal{X}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}$ requires the data of lifts of $\mathbbm{b}$, $\mathbbm{b}^{\prime}$ into PL functions on $\bf M$$_{\mathbb{R}}$ by Remark 3.39.
Proof.
Let $\mathcal{L}$} and $\mathcal{L}$$^{\prime}$ be the defining relatively ample line bundles of the two Mumford constructions over the analytic tubular neighborhood $T({\mathbb{B}})\subset Y({\mathbb{B}})$ of the deepest toric stratum.
Take a $\bf M$}-periodic, regular refinement $\mathcal{S}$} of the normal fans for $\mathbbm{b}$, $\mathbbm{b}^{\prime}$, and let $X^{\rm univ}_{\circ}({\mathcal{S}})\to T({\mathbb{B}})$ be the corresponding fan Construction 3.11. Let ${\mathcal{L}}_{\mathcal{S}}$ and ${\mathcal{L}}_{\mathcal{S}}^{\prime}$ be the pullbacks of $\mathcal{L}$} and $\mathcal{L}$$^{\prime}$ to $X^{\rm univ}({\mathcal{S}})$, and define ${\mathcal{E}}:={\mathcal{L}}_{\mathcal{S}}^{\prime}\otimes{\mathcal{L}}_{\mathcal{S}}^{-1}$. Finally, let $\widetilde{{\mathcal{E}}}$, $\widetilde{{\mathcal{L}}}_{\mathcal{S}}$, $\widetilde{{\mathcal{L}}}\,\!^{\prime}_{S}$ be the pullbacks of $\mathcal{E}$}, ${\mathcal{L}}_{\mathcal{S}}$, ${\mathcal{L}}_{{\mathcal{S}}}^{\prime}$ to the universal cover of $X^{\rm univ}_{\circ}({\mathcal{S}})$. Then $\widetilde{{\mathcal{E}}}$, $\widetilde{{\mathcal{L}}}_{\mathcal{S}}$, $\widetilde{{\mathcal{L}}}\,\!^{\prime}_{S}$ are $\bf M$}-equivariant line bundles. The condition that $b^{\prime}-b$ is $\bf M$}-periodic implies that we have an $\bf M$}-equivariant isomorphism
$\widetilde{{\mathcal{E}}}\simeq{\mathcal{O}}(\sum_{\begin{subarray}{c}\textrm{rays}\\ {\mathbb{R}}_{\geq 0}v_{i}\in{\mathcal{S}}\end{subarray}}a_{v_{i}}D_{v_{i}}),$
with $a_{v_{i}}\in{\mathbb{Z}}$ depending only on the $\bf M$}-equivalence class $\overline{v}_{i}$ of the ray ${\mathbb{R}}_{\geq 0}v_{i}\in{\mathcal{S}}$. Quotienting, we deduce that the line bundle $\mathcal{E}$} is represented by a finite $\mathbb{Z}$}-linear sum $\sum a_{\overline{v}_{i}}D_{\overline{v}_{i}}$ of components over the boundary of $X^{\rm univ}_{\circ}({\mathcal{S}})$. As components over the boundary, $D_{\overline{v}_{i}}$ descend to algebraic divisors on the algebraic space $X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ and thus ${\mathcal{L}}_{\mathcal{S}}$ and ${\mathcal{L}}^{\prime}_{{\mathcal{S}}}$ differ by twisting by a linear combination of vertical divisors, over the boundary of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. So one extends as a section of relative Picard if and only if the other does. Furthermore, the relative ampleness of $\mathcal{L}$} and $\mathcal{L}$$^{\prime}$ over the interior $\widetilde{{\mathcal{A}}}_{g}$ are equivalent. On the other hand, the relative ampleness of either over the boundary $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}\setminus\widetilde{{\mathcal{A}}}_{g}$ is automatic, by construction. ∎
Proposition 6.16.
Let $f\colon X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ be a universal polytopal Mumford construction associated to a lift of $\mathbbm{b}\subset H^{0}({\mathbb{T}}^{g},\tfrac{1}{d}{\mathbb{Z}}{\rm PL}/\tfrac{1}{d}{\mathbb{Z}}{\rm L})$, extending (as an algebraic space) the universal family $\widetilde{{\mathcal{X}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}$ of abelian varieties. Then $f$ is a projective morphism whenever $b_{i}({\bf m})-b_{i}(-{\bf m})$ is $\bf M$}-periodic. Similarly, there is a relatively projective extension $f\colon X^{{\rm univ}\star}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}$ of the universal family $\widetilde{{\mathcal{X}}}_{g}^{\star}\to\widetilde{{\mathcal{A}}}_{g}$ of abelian torsors when $\overline{b}\sim\overline{b}\,\!^{\rm vor}$ lie in the same shift class, for any $B\in{\mathbb{B}}$, and for $\overline{b}\,\!^{\rm vor}$ defined as in (36).
Proof.
To prove the extension result for $\widetilde{{\mathcal{X}}}_{g}$, we follow the proof strategy of Proposition 6.15: We may pass to a smooth (1)(-1)-symmetric common refinement $X^{\rm univ}_{\circ}({\mathcal{S}})\to T({\mathbb{B}})$ of the Mumford constructions for $b_{i}({\bf m})$ and $b_{i}(-{\bf m})$. The pulled back line bundles ${\mathcal{L}}_{S}$, ${\mathcal{L}}_{\mathcal{S}}^{\prime}$ associated to $b_{i}({\bf m})$, $b_{i}(-{\bf m})$ are interchanged by the (1)(-1)-involution: $(-1)^{*}{\mathcal{L}}_{\mathcal{S}}\simeq{\mathcal{L}}_{{\mathcal{S}}}^{\prime}$. If $b_{i}({\bf m})-b_{i}(-{\bf m})$ is $\bf M$}-periodic, we may conclude that ${\mathcal{L}}_{{\mathcal{S}}}^{\prime}\simeq{\mathcal{L}}_{{\mathcal{S}}}(\,\sum a_{\overline{v}_{i}}D_{\overline{v}_{i}})$ differ by twisting by vertical divisors. It follows that $({\mathcal{L}}_{\mathcal{S}})^{\otimes 2}$ defines an algebraic extension of $({\mathcal{L}}\,\!_{\widetilde{{\mathcal{X}}}_{g}})^{\otimes d}$ where ${\mathcal{L}}\,\!_{\widetilde{{\mathcal{X}}}_{g}}\in{\rm Pic}_{\widetilde{{\mathcal{X}}}_{g}/\widetilde{{\mathcal{A}}}_{g}}(\widetilde{{\mathcal{A}}}_{g})$ is a tensor square of a (1)(-1)-symmetric local lift of the principal polarization.
The case of extending $\widetilde{{\mathcal{X}}}_{g}^{\star}$ is similar, but again Proposition 6.15 does not directly apply, since we are gluing onto the universal abelian torsor. The extension of $\widetilde{{\mathcal{X}}}_{g}^{\star}$ by the Mumford construction of second Voronoi type (Construction 6.6) is relatively projective. Replacing $\mathbb{B}$} with a common refinement of $\mathbb{B}$} and ${\mathfrak{F}}_{\rm vor}$, we may assume that $\mathbb{B}$} is contained in a second Voronoi cone. Choosing lifts $b^{\rm vor}$ as in (36), defines a local analytic section of $\overline{{\mathcal{X}}_{g}^{\star}}^{\rm vor}$ near the boundary stratum associated to a second Voronoi cone. The hypothesis that $\overline{b}\sim\overline{b}\,\!^{\rm vor}$, for all $B\in{\mathbb{B}}$ and argument of Proposition 6.15 show that, with respect to the chosen local origin section of $\widetilde{{\mathcal{X}}}_{g}^{\star}$, and a well chosen lift of $\mathbbm{b}$, there is a gluing $X^{{\rm univ}\star}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ of $X^{{\rm univ}}(\mathbbm{b})\to T({\mathbb{B}})$ and $\widetilde{{\mathcal{X}}}_{g}^{\star}\to\widetilde{{\mathcal{A}}}_{g}$ for which the ample line bundle on the former extends to a section in ${\rm Pic}_{X^{{\rm univ}\star}(\mathbbm{b})/\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}}(\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}})$.
Thus, in either case $\widetilde{{\mathcal{X}}}_{g}$ or $\widetilde{{\mathcal{X}}}_{g}^{\star}$, the canonical polarization on the Mumford construction extends to a relatively ample section of relative Picard over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$.
After passing to some further tensor power, we may lift to an element $\mathcal{E}$} of the (algebraic) Picard group of $X^{\rm univ}(\mathbbm{b})$ or $X^{{\rm univ}\star}(\mathbbm{b})$, which is relatively very ample. Pushing forward, we get a vector bundle $f_{*}{\mathcal{E}}$ over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ over the étale site, and therefore over the Zariski site. It follows that the projectivization of $(f_{*}{\mathcal{E}})^{\vee}$ is relatively projective over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. Thus, $X^{\rm univ}(\mathbbm{b})$ or $X^{{\rm univ}\star}(\mathbbm{b})$ admit closed, algebraic embeddings into a projective space over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. The result follows. ∎
Projectivity criteria for extensions of $\widetilde{{\mathcal{X}}}_{g}$ should be compared to [17, Ch. VI]. Following a standard toric construction, the polytope $\Gamma$ of Section 3.3 defines a convex PL function $\tilde{\phi}\colon{\mathcal{S}}\to{\mathbb{R}}$ on the normal fan $\mathcal{S}$} of Section 3.1. The function $\tilde{\phi}$ should be an “admissible homogeneous principal polarization function” as in [17, Ch. VI, Def. 1.5], with our conditions on the shift class of $b_{i}$ related to Def. 1.5.(vi) of loc.cit. Our projectivity results should then follow from [17, Ch. VI, Thm. 1.13], though translating between the language used here and that in loc.cit. is somewhat involved.
Corollary 6.17.
Suppose that $\mathscr{H}$ is a hyperplane arrangement for the regular matroid $\underline{R}$}, for which the parallel hyperplanes normal to $\vec{x}_{i}$ are $H_{i}^{(j)}\coloneqq\vec{x}_{i}({\bf m})\in\epsilon_{ij}+{\mathbb{Z}}$. Then the extension
$X^{{\rm univ}\star}({\underline{R}},\mathscr{H})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{{\underline{R}}}}$
of $\widetilde{{\mathcal{X}}}_{g}^{\star}\to\widetilde{{\mathcal{A}}}_{g}$ is projective whenever $\sum_{j}\epsilon_{ij}=0\in{\mathbb{Q}}/{\mathbb{Z}}$ for all $i=1,\dots,k$ and the extension
$X^{\rm univ}({\underline{R}},\mathscr{H})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{{\underline{R}}}}$
of $\widetilde{{\mathcal{X}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}$ is projective whenever $\sum_{j}(\epsilon_{ij}+\tfrac{1}{2})=0\in{\mathbb{Q}}/{\mathbb{Z}}$ for all $i=1,\dots,k$.
Proof.
Let $r_{i}$ be the number of hyperplanes $H_{i}^{(j)}\in\mathscr{H}$ normal to $\vec{x}_{i}$.
For ${\mathbb{B}}_{\underline{R}}\in{\mathfrak{F}}_{\rm vor}$ a matroidal cone, the Delaunay decomposition as in Construction 6.5 is the unshifted hyperplane arrangement (28). Thus, it follows from Proposition 6.16 that $\mathscr{H}$ defines a projective extension of $\widetilde{{\mathcal{X}}}_{g}^{\star}\to\widetilde{{\mathcal{A}}}_{g}$ whenever the section $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},\tfrac{1}{d}{\mathbb{Z}}{\rm PL}/\tfrac{1}{d}{\mathbb{Z}}{\rm L})$ bending along $H_{i}^{(j)}:=\{\vec{x}_{i}({\bf m})\in\epsilon_{ij}+{\mathbb{Z}}\}$ lies in the same shift class as $\overline{b}_{i}^{o}\in H^{0}({\mathbb{T}}^{g},\tfrac{1}{d}{\mathbb{Z}}{\rm PL}/\tfrac{1}{d}{\mathbb{Z}}{\rm L})$ which bends $r_{i}$ times along the unshifted hyperplane $\vec{x}_{i}({\bf m})\in 0+{\mathbb{Z}}$. Equivalently,
(37) $\textstyle\int_{0}^{1}\int_{0^{-}}^{x}(-r_{i}\delta_{0}+\sum_{j=1}^{r_{i}}\delta_{\epsilon_{ij}})\,dy\,dx\in{\mathbb{Z}}$
where $\delta_{p}$ denotes the Dirac delta function at $p$. Now (37) holds if and only if $\sum_{j=1}^{r_{i}}\epsilon_{ij}\in{\mathbb{Z}}$.
For the case of extending $\widetilde{{\mathcal{X}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}$ projectively, we observe that an arrangement which bends $r_{i}$ times along the half-shifted hyperplane $\vec{x}_{i}({\bf m})\in\tfrac{1}{2}+{\mathbb{Z}}$ satisfies the hypotheses of Proposition 6.16 by lifting to a $\tfrac{1}{2}{\mathbb{Z}}{\rm PL}$ function $b_{i}^{\prime}$ which is identically zero in a neighborhood of the origin of $\bf M$$_{\mathbb{R}}$. So $b_{i}^{\prime}$ defines a relatively projective extension of the universal family $\widetilde{{\mathcal{X}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}$. Then by Proposition 6.15, it suffices to understand when $\overline{b}_{i}^{\prime}\sim\overline{b}_{i}$ lie in the same shift class, i.e.
$\textstyle\int_{0}^{1}\int_{0^{-}}^{x}(-r_{i}\delta_{\frac{1}{2}}+\sum_{j=1}^{r_{i}}\delta_{\epsilon_{ij}})\,dy\,dx\in{\mathbb{Z}}.$
This holds exactly when $\sum_{j=1}^{r_{i}}(\epsilon_{ij}+\tfrac{1}{2})\in{\mathbb{Z}}$. ∎
@@
Our goal is to prove Theorem 1.1, and leverage Proposition 6.12 to prove more algebraic formulations of the results therein, see Theorem 7.1 and Corollary 7.2 below.
@@
Let $f^{\ast}\colon X^{\ast}\to(\Delta^{*})^{k}$ be a family of PPAVs which is matroidal with respect to the snc extension $(\Delta^{*})^{k}\hookrightarrow\Delta^{k}$ (Def. 4.9). Then there are integers $r_{i}>0$ for which the monodromy bilinear forms about $\{u_{i}=0\}$ are $r_{i}B_{i}$ where $B_{i}={\bf x}_{i}^{2}$ is an integral generator of the matroidal cone ${\mathbb{B}}_{\underline{R}}$ of the corresponding regular matroid $\underline{R}$}. Constructions 4.16 and 3.26, see also Notation 4.19, give a universal Mumford degeneration
@@
The monodromies about the boundary divisors of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$ are exactly $r_{i}B_{i}$ and by Proposition 2.27, the classifying morphism $(\Delta^{*})^{k}\to{\mathcal{A}}_{g}$ will (lift and) extend to $\Delta^{k}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$. Pulling back $X^{\rm univ}({\underline{R}},\mathscr{H})$ along the extension of the classifying morphism to $\Delta^{k}$ produces the desired extension $f\colon X\to\Delta^{k}$. It has smooth total space and nodal singularities, because $f^{\rm univ}$ is locally trivial along the deepest toroidal stratum of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$ and so smoothness, resp. nodality, of $X^{\rm univ}$, resp. $f^{\rm univ}$ (see Proposition 5.3), implies smoothness, resp. nodality, of the restrictions $X$, resp. $f$, to the transversal slice $\Delta^{k}$ to this deepest toroidal stratum.
@@
The second part of Theorem 1.1 follows from the existence of algebraic, transversely shifted matroidal degenerations with specified monodromies, see Corollary 7.2 below. ∎
@@
Let $f^{\ast}\colon X^{\ast}\to Y^{\ast}$ be a projective family of PPAVs over a base $Y^{*}=Y\setminus D$ for $D\subset Y$ an snc divisor in a smooth quasiprojective variety $Y$. Let $0\in D$ and assume the local monodromy bilinear forms $B_{i}$ about the components $D_{i}\ni 0$ are $r_{i}{\bf x}_{i}^{2}$ for an integral realization $i\mapsto{\bf x}_{i}\in{\bf M}^{\vee}$ of a regular matroid $\underline{R}$}, where ${\bf M}\simeq{\rm gr}^{W}_{0}H_{1}(X_{t},{\mathbb{Z}})$ for $t$ near 0. Up to passing to an étale neighborhood of $0\in Y$, there is a flat, projective, $D$-nodal, relatively $K$-trivial extension $f\colon X\to Y$, which is furthermore strictly $D$-nodal when all $r_{i}\geq 2$.